You recently purchased a stock that is expected to earn 12 percent in a booming economy, 9 percent in a normal economy and lose 4 percent in a recessionary economy. There is a 13 percent probability of a boom, a 75 percent chance of a normal economy. What is your expected rate of return on this stock? 8.79 8.18 4.00 7.18 7.83
Answer #1

Probability of recession=(100-13-75)=12%

Expected rate of return=Respective return*Respective Probability

=(12*0.13)+(9*0.75)-(4*0.12)

=7.83%